OpenStudy (anonymous):

simplify first. then solve using logs. 1322.17/180000=(.03897/12)/(1-(1+(.03897/12))^(-12t))

OpenStudy (anonymous):

wowwwwwwwwwwwwwwwwww i am soooo not in your grade

OpenStudy (anonymous):

kayalakizz i don't think you you'll want to try to do this.

OpenStudy (anonymous):

nope not at all

OpenStudy (anonymous):

go on chat

OpenStudy (anonymous):

\[\frac{1322.17}{180000}=\frac{(\frac{.03897}{12})}{(1-(1+(\frac{.03897}{12}))^{-12t})}\]